# Unified Privacy Guarantees for Decentralized Learning via Matrix Factorization

### v1

@@ 1 Introduction
In Decentralized Learning (DL), participants collaboratively train a shared model by exchanging model parameters on a peer-to-peer communication graph (Lian et al., 2017, 2018; Koloskova et al., 2020; Beltrán et al., 2022; Tian et al., 2023; Yuan et al., 2024). By removing the central server, DL offers faster deployment, higher scalability, and improved robustness. Since data remains local and each participant sees only a subset of communications, DL is also motivated by privacy concerns (Cyffers & Bellet, 2022; Cyffers et al., 2022; Kairouz et al., 2021c). However, decentralization alone is insufficient to ensure privacy, as exchanged messages can still leak sensitive information that enables inference or reconstruction of local data (El Mrini et al., 2024; Pasquini et al., 2023). Data protection thus requires additional mechanisms beyond decentralization.
Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for privacy in machine learning. DP protects participation by ensuring that outputs do not depend too strongly on any single data point, typically through noise injection. Extending DP to Decentralized Learning is challenging: unlike in centralized settings, where a trusted curator releases only the final output, DL exposes intermediate peer-to-peer messages to participants. A natural option is to analyze DP-DL algorithms under Local DP (Kasiviswanathan et al., 2008; Bellet et al., 2018; Wang et al., 2015), which assumes all messages are released publicly, but this overly conservative trust model often leads to poor privacy–utility trade-offs (Chan et al., 2012). More realistic models, such as Pairwise Network DP (PNDP) (Cyffers et al., 2022; Cyffers & Bellet, 2022) and Secret-based LDP (SecLDP) (Allouah et al., 2024), have recently been proposed to account for adversaries with only partial knowledge of the system, showing that decentralization can amplify privacy guarantees relative to Local DP.
@@ 2 Related work
Matrix factorization mechanism. The Matrix Factorization (MF) mechanism for SGD first appeared as a generalization of correlated noise in the online setting (Kairouz et al., 2021b), based on tree mechanisms (Dwork et al., 2010; Jain et al., 2012), and was later formalized by Denisov et al. (2022). More recent work improves the factorization either by designing more efficient decompositions (Choquette-Choo et al., 2023), or by allowing additional constraints on the strategy matrix, i.e., the matrix encoding locally applied noise correlations on the gradient. Existing constraints typically enforce a limited window for temporal correlations via banded matrices, which reduces computation (Choquette-Choo et al., 2023; Kalinin & Lampert, 2024), whereas in this work our constraint in Mafalda-SGD enforces that correlations occur only within nodes. Other works extend the mechanism to more complex participation schemes (Choquette-Choo et al., 2023). For a more comprehensive survey, we refer the reader to Pillutla et al. (2025).
Differentially private decentralized learning. Private decentralized learning was first studied under the trust model of Local DP (Bellet et al., 2018; Cheng et al., 2019; Huang et al., 2015), which suffers from poor utility due to overly pessimistic assumptions: participants are only protected by their own noise injection, since all messages are assumed to be known to the attacker. Relaxed trust models have since gained traction, such as Pairwise Network Differential Privacy (PNDP) (Cyffers et al., 2022), initially introduced for gossip averaging and later extended to several algorithms (Cyffers et al., 2024; Li et al., 2025; Biswas et al., 2025). PNDP assumes that the attacker is a node participating in the algorithm, and that it is thus sufficient to protect what a node observes during the execution—typically only the messages it sends or receives. Similarly, SecLDP (Allouah et al., 2024) enforces DP conditional on a set of secrets that remain hidden from the attacker. The idea of correlating noise has a long history in DL, first as an obfuscation method against neighbors without formal guarantees (Mo & Murray, 2017; Gupta et al., 2020). More recent works provide DP guarantees for algorithms with correlated noise (Sabater et al., 2022; Allouah et al., 2024; Biswas et al., 2025), but correlations are not optimized and instead dictated by what can be proven, which explains why our method Mafalda-SGD significantly outperforms them. Finally, the concept of representing attacker knowledge in DL via a matrix that can be constructed algorithmically was introduced by El Mrini et al. (2024) and more recently studied by Koskela & Kulkarni (2025), but only for specific cases of algorithms and trust models, and without establishing a connection to the matrix factorization mechanism.
@@ 3 Background: Matrix Factorization Mechanism
For a fixed sensitivity $\operatorname*{sens}_{\Pi}(C)$, generating $Z\sim\mathcal{N}(0,\sigma^{2}\operatorname*{sens}_{\Pi}(C)^{2})^{T\times d}$ ensures that $(G+C^{\dagger}Z)$ is $\tfrac{1}{\sigma}$-GDP. Optimizing the noise correlation is typically done by minimizing $\operatorname*{sens}(C)^{2}\left\lVert B\right\rVert^{2}$ under the constraint $A^{\text{pre}}=BC$, where $\left\lVert B\right\rVert^{2}$ represents the utility of the factorization. While the privacy analysis is straightforward if $G,G^{\prime}$ are fixed in advance, the guarantee remains valid when $G$ is computed adaptively through time, with $g_{t}$ depending on $g_{1},\dots,g_{t-1}$. This has been proven only for workload matrices that are square, lower triangular, and full rank (Denisov et al., 2022; Pillutla et al., 2025). We show in Section 5 that the privacy guarantees also hold in a more general case, thereby enabling adaptivity in DL as well.
@@ 4.1 Warm-up: MF-D-SGD under Local DP
Model updates. We stack the local parameters of the $n$ nodes as a vector $\theta_{t}\in\mathbb{R}^{n\times d}$, where $\theta_{t}^{(u)}$ is the local parameter of node $u$ at time $t\in\left\llbracket 0,T-1\right\rrbracket$. The gradient step can then be written as $\theta_{t+\frac{1}{2}}=\theta_{t}-\eta\,\bigl(G_{t}+C^{\dagger}_{t}Z\bigr),$ where each row $u$ of $G_{t}$ is the vector $G_{t}^{(u)}=\nabla f_{u}\bigl(\theta_{t}^{(u)},\xi_{t}^{(u)}\bigr)^{\top}$, the gradient of the local loss function $f_{u}$ on the sample $\xi_{t}^{(u)}$ from the local dataset $\mathcal{D}_{u}$, and $C^{\dagger}_{t}\in\mathbb{R}^{n\times m}$ is the block of $C^{\dagger}$ from the $(tn+1)$-th row to the $(t+1)n$-th one. Here, $C^{\dagger}_{t}Z$ can lead to arbitrarily complex noise correlation. Note that for $C=I_{nT}$, we recover the local updates of DP-D-SGD, the simple decentralized SGD with uncorrelated noise. After the local updates, nodes perform a gossiping step, exchanging their models with neighbors and averaging the intermediate models $\theta_{t+\tfrac{1}{2}}$ using the gossip matrix $W$.
@@ 5 More General Privacy Guarantees for Matrix Factorization
Let $B$ and $C$ be two matrices. The sensitivity of $C$ with respect to $B$ and a participation schema $\Pi$ is
$\operatorname*{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\left\|C\left(G-G^{\prime}\right)\right\|_{B^{\dagger}B}.$ (7)
@@
Theorem 7 shows that the privacy guarantees of a matrix factorization mechanism remain valid for a broader class of matrices $A$, even when the gradient vector is chosen adaptively, i.e., when it may depend arbitrarily on all past information (Denisov et al., 2022). The column-echelon form property of the matrix ensures it cannot depend on future information.
Our result generalizes existing results by allowing $A$ to be rectangular and possibly rank-deficient, a relaxation crucial for applying it to DL algorithms. The price of this extension is a dependency on $B$: since $C$ is not invertible, $B$ is not unique, and sensitivity now depends on both the encoder matrix $C$ and the decoder matrix $B$. The correction $B^{\dagger}B$ can be seen as a projection onto the column space of $B$, discarding inaccessible gradient combinations. When $B$ is square and full rank, as in prior work, we recover the known formula $\operatorname*{sens}_{\Pi}(C;B)=\operatorname*{sens}_{\Pi}(C;I)=\operatorname*{sens}_{\Pi}(C)$.
@@
In DL, the participation schema $\Pi$ is also slightly modified. A participation pattern $\pi$ is associated to a local dataset ${\cal D}_{u}$ and can be non-negative only for the rows in $\{u,n+u,\dots,(T-1)n+u\}$, which corresponds to gradients computed from ${\cal D}_{u}$.
@@ 6 A Novel Algorithm with Optimized Noise Correlation
where $L$ is lower triangular and obtained by Cholesky decomposition of $H$.
This yields a standard factorization problem for $L$, which can be solved using existing frameworks for the MF mechanism Choquette-Choo et al. (2023); Pillutla et al. (2025). In practice, these frameworks operate on the Gram matrix $H=L^{\top}L$, avoiding the need for Cholesky decomposition.
@@
5:return MF-D-SGD($W,C_{\text{mafalda}},T,\Delta_{g},\sigma,{\theta}_{0}$) $\triangleright$ Algorithm 1
@@ 7.1 Tighter accounting for PNDP
  • Facebook Ego graph (Leskovec & Mcauley, 2012): corresponds to the graph of a fixed Facebook user (omitted from the graph), where nodes are the user’s friends and edges represent friendships between them. This graph has $148$ nodes.
  • PeerTube graph (Damie & Cyffers, 2025): an example of decentralized social platform. PeerTube is an open-source decentralized alternative to YouTube. Each node represents a PeerTube server and edges encode follow relationships between servers, allowing users to watch recommended videos across instances. The graph, restricted to its largest connected component, has $271$ nodes.
  • Florentine Families (Breiger & Pattison, 1986): a historical graph with $15$ nodes describing marital relations between families in 15th-century Florence.
@@ 7.2 Mafalda-SGD
We study a regression task on the Housing dataset, [^1] https://www.openml.org/d/823 which has $8$ features and 20,64020{,}640 data points, with the goal of predicting house prices based on demographic and housing features. We use a simple multilayer perceptron (MLP) with one hidden layer of width $64$ and ReLU activation, followed by a linear output layer for regression, trained with the mean-square error loss, in all experiments. This dataset has been used in several related works (Cyffers et al., 2022; Li et al., 2025). In all experiments, we cycle over the data points such that each of them participates $20$ times in total, with participation every $19$ steps. These numbers were chosen for memory constraints and not optimized. This cyclic participation corresponds to a $(k,b)$-participation scheme with $k=20$ and $b=19$ in the matrix factorization literature (Choquette-Choo et al., 2023).
@@ Acknowledgments
This work was supported by grant ANR-20-CE23-0015 (Project PRIDE) and the ANR 22-PECY-0002 IPOP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR and by the Austrian Science Fund (FWF) [10.55776/COE12]. This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011015352R1 made by GENCI. The authors thank Quino for his comic strip Mafalda which inspired our new algorithm name.

### v2

@@ 1 Introduction
In Decentralized Learning (DL), participants collaboratively train a shared model by exchanging model parameters on a peer-to-peer communication graph (Lian et al., 2017; 2018; Koloskova et al., 2020; Beltrán et al., 2022; Tian et al., 2023; Yuan et al., 2024). By removing the central server, DL offers faster deployment, higher scalability, and improved robustness. Since data remains local and each participant sees only a subset of communications, DL is also motivated by privacy concerns (Cyffers and Bellet, 2022; Cyffers et al., 2022; Kairouz et al., 2021c). However, decentralization alone is insufficient to ensure privacy, as exchanged messages can still leak sensitive information that enables inference or reconstruction of local data (El Mrini et al., 2024; Pasquini et al., 2023). Data protection thus requires additional mechanisms beyond decentralization.
Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for privacy in machine learning. DP protects participation by ensuring that outputs do not depend too strongly on any single data point, typically through noise injection. Extending DP to Decentralized Learning is challenging: unlike in centralized settings, where a trusted curator releases only the final output, DL exposes intermediate peer-to-peer messages to participants. A natural option is to analyze DP-DL algorithms under Local DP (Kasiviswanathan et al., 2008; Bellet et al., 2018; Wang et al., 2015), which assumes all messages are released publicly, but this overly conservative trust model often leads to poor privacy–utility trade-offs (Chan et al., 2012). More realistic models, such as Pairwise Network DP (PNDP) (Cyffers et al., 2022; Cyffers and Bellet, 2022) and Secret-based LDP (SecLDP) (Allouah et al., 2024), have recently been proposed to account for adversaries with only partial knowledge of the system, showing that decentralization can amplify privacy guarantees relative to Local DP.
@@ 2 Related work
Matrix factorization mechanism. The Matrix Factorization (MF) mechanism for SGD first appeared as a generalization of correlated noise in the online setting (Kairouz et al., 2021b), based on tree mechanisms (Dwork et al., 2010; Jain et al., 2012), and was later formalized by Denisov et al. (2022). More recent work improves the factorization either by designing more efficient decompositions (Choquette-Choo et al., 2023a), or by allowing additional constraints on the strategy matrix, i.e., the matrix encoding locally applied noise correlations on the gradient. Existing constraints typically enforce a limited window for temporal correlations via banded matrices, which reduces computation (Choquette-Choo et al., 2023b; Kalinin and Lampert, 2024), whereas in this work our constraint in Mafalda-SGD enforces that correlations occur only within nodes. Other works extend the mechanism to more complex participation schemes (Choquette-Choo et al., 2023c). For a more comprehensive survey, we refer the reader to Pillutla et al. (2025).
Differentially private decentralized learning. Private decentralized learning was first studied under the trust model of Local DP (Bellet et al., 2018; Cheng et al., 2019; Huang et al., 2015), which suffers from poor utility due to overly pessimistic assumptions: participants are only protected by their own noise injection, since all messages are assumed to be known to the attacker. Relaxed trust models have since gained traction, such as Pairwise Network Differential Privacy (PNDP) (Cyffers et al., 2022), initially introduced for gossip averaging and later extended to several algorithms (Cyffers et al., 2024; Li et al., 2025; Biswas et al., 2025). PNDP assumes that the attacker is a node participating in the algorithm, and that it is thus sufficient to protect what a node observes during the execution—typically only the messages it sends or receives. Similarly, SecLDP (Allouah et al., 2024) enforces DP conditional on a set of secrets that remain hidden from the attacker. The idea of correlating noise has a long history in DL, first as an obfuscation method against neighbors without formal guarantees (Mo and Murray, 2017; Gupta et al., 2020). More recent works provide DP guarantees for algorithms with correlated noise (Sabater et al., 2022; Allouah et al., 2024; Biswas et al., 2025), but correlations are not optimized and instead dictated by what can be proven, which explains why our method Mafalda-SGD significantly outperforms them. Finally, the concept of representing attacker knowledge in DL via a matrix that can be constructed algorithmically was introduced by El Mrini et al. (2024) and more recently studied by Koskela and Kulkarni (2025), but only for specific cases of algorithms and trust models, and without establishing a connection to the matrix factorization mechanism.
@@ 3 Background: Matrix Factorization Mechanism
For a fixed sensitivity $\operatorname*{sens}_{\Pi}(C)$, generating $Z\sim\mathcal{N}(0,\sigma^{2}\operatorname*{sens}_{\Pi}(C)^{2})^{T\times d}$ ensures that $(G+C^{\dagger}Z)$ is $\tfrac{1}{\sigma}$-GDP. Optimizing the noise correlation is typically done by minimizing $\operatorname*{sens}(C)^{2}\left\lVert B\right\rVert^{2}$ under the constraint $A^{\text{pre}}=BC$, where $\left\lVert B\right\rVert^{2}$ represents the utility of the factorization. While the privacy analysis is straightforward if $G$ and $G^{\prime}$ are fixed in advance, the guarantee remains valid when $G$ is computed adaptively through time, with $g_{t}$ depending on $g_{1},\dots,g_{t-1}$. This has been proven only for workload matrices that are square, lower triangular, and full rank (Denisov et al., 2022; Pillutla et al., 2025). We show in Section 5 that the privacy guarantees also hold in a more general case, thereby enabling adaptivity in DL as well.
@@ 4.1 Warm-up: MF-D-SGD under Local DP
Model updates. We stack the local parameters of the $n$ nodes as a vector $\theta_{t}\in\mathbb{R}^{n\times d}$, where $\theta_{t}^{(u)}$ is the local parameter of node $u$ at time $t\in\left\llbracket 0,T-1\right\rrbracket$. The gradient step can then be written as $\theta_{t+\frac{1}{2}}=\theta_{t}-\eta\,\bigl(G_{t}+C^{\dagger}_{t}Z\bigr),$ where each row $u$ of $G_{t}$ is the vector $G_{t}^{(u)}=\nabla f_{u}\bigl(\theta_{t}^{(u)},\xi_{t}^{(u)}\bigr)^{\top}$, and $C^{\dagger}_{t}\in\mathbb{R}^{n\times m}$ is the block of $C^{\dagger}$ from the $(tn+1)$-th row to the $(t+1)n$-th one. $C^{\dagger}_{t}Z$ captures arbitrary noise correlation, and setting $C=I_{nT}$ recovers the local updates of DP-D-SGD, the simple decentralized SGD with uncorrelated noise. After the local updates, nodes perform a gossiping step, exchanging their models with neighbors and averaging the intermediate models $\theta_{t+\tfrac{1}{2}}$ using the gossip matrix $W$.
@@
Remark 3.
Up to this point, we have assumed that the gossip matrix $W$ is fixed across iterations for simplicity’s sake (notably when explaining Figure 1). In the remainder of the paper we move to the more general case of time-varying gossip matrices $W_{t}$. This change only requires replacing powers of a single matrix gossip $W$ by products of matrices in the expression of ${\mathbf{W}}_{T}$. See Section A.7 for details of this more general expression.
@@ 5 More General Privacy Guarantees for Matrix Factorization
Let $B$ and $C$ be two matrices. The sensitivity of $C$ with respect to $B$ and a participation scheme $\Pi$ is
$\operatorname*{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\left\|C\left(G-G^{\prime}\right)\right\|_{B^{\dagger}B},$ (7)
with $\left\lVert A\right\rVert_{B}^{2}:=\operatorname*{tr}\left(A^{\top}BA\right)$.
@@
Theorem 8 shows that the privacy guarantees of a matrix factorization mechanism remain valid for a broader class of matrices $A$, even when the gradient vector is chosen adaptively, i.e., when it may depend arbitrarily on all past information (Denisov et al., 2022). The column-echelon form property of the matrix ensures it cannot depend on future information. This generalization of the usual lower-triangular property is necessary when considering attackers who have only partial information over the network. Most sets of observations that follow the natural causal ordering in DL can be rewritten as a column-echelon matrix.
Our result generalizes existing results by allowing $A$ to be rectangular and possibly rank-deficient, a relaxation crucial for applying it to DL algorithms. The price of this extension is a dependency on $B$: since $B$ may not be invertible, $C$ is not unique, and the sensitivity now depends on both the encoder matrix $C$ and the decoder matrix $B$. The correction $B^{\dagger}B$ can be interpreted as a projection of the columns of $C$ onto the row space of $B$, discarding unobserved gradient combinations. When $B$ is square and full rank, as in prior work, we recover the known formula $\operatorname*{sens}_{\Pi}(C;B)=\operatorname*{sens}_{\Pi}(C;I)=\operatorname*{sens}_{\Pi}(C)$.
@@
In DL, the participation scheme $\Pi$ is also slightly modified. A participation pattern $\pi$ is associated to a local dataset ${\cal D}_{u}$ and can be non-negative only for the rows in $\{u,n+u,\dots,(T-1)n+u\}$, which corresponds to gradients computed from ${\cal D}_{u}$.
Remark 10.
Following Corollary 1 of Dong et al. (2022), $\mu$-GDP directly and tightly translates into $(\epsilon,\delta)$-DP, with appropriate $\epsilon$ and $\delta$ values. Moreover, it is also possible to derive $(\alpha,\epsilon)$-Rényi DP guarantees Gil et al. (2013).
@@ 6 A Novel Algorithm with Optimized Noise Correlation
where $L$ is lower triangular and obtained by Cholesky decomposition of $H=L^{\top}L$.
This yields a standard factorization problem for $L$, which can be solved using existing frameworks for the MF mechanism Choquette-Choo et al. (2023c); Pillutla et al. (2025). In practice, these frameworks operate on the Gram matrix $H=L^{\top}L$, avoiding the need for Cholesky decomposition.
@@
[…]
6:return MF-D-SGD($W,C_{\text{mafalda}},T,\Delta_{g},\sigma,{\theta}_{0}$) $\triangleright$ Algorithm 1
@@
Remark 13.
The complexity of Algorithm 2 is driven by (i) the computation of the matrix $L$, which costs $\mathcal{O}(n^{3}T+n^{2}T^{3})$ in time, and $\mathcal{O}(n^{2}T^{2})$ in space, and (ii) the minimization of $\mathcal{L}_{\text{opti}}$ which is $\mathcal{O}(T^{3})$ per step of L-BFGS. Correlation restarts (Pillutla et al., 2025) can be used to keep $T$ within reasonable bounds.
@@ 7.1 Tighter accounting for PNDP
  • Facebook Ego graph (Leskovec and Mcauley, 2012): corresponds to the graph of a fixed Facebook user (omitted from the graph), where nodes are the user’s friends and edges represent friendships between them. This graph has $148$ nodes.
  • PeerTube graph (Damie and Cyffers, 2026): an example of decentralized social platform. PeerTube is an open-source decentralized alternative to YouTube. Each node represents a PeerTube server and edges encode follow relationships between servers, allowing users to watch recommended videos across instances. The graph, restricted to its largest connected component, has $271$ nodes.
  • Florentine Families (Breiger and Pattison, 1986): a historical graph with $15$ nodes describing marital relations between families in 15th-century Florence.
@@ 7.2 Mafalda-SGD
We study a regression task on the Housing dataset, [^1] https://www.openml.org/d/823 which has $8$ features and 20,64020{,}640 data points, with the goal of predicting house prices based on demographic and housing features. In all experiments, we use a simple multilayer perceptron (MLP) with one hidden layer of width $64$ and ReLU activation, followed by a linear output layer for regression, trained with the mean-square error loss. This dataset has been used in several related works (Cyffers et al., 2022; Li et al., 2025). In all experiments, we cycle over the data points such that each of them participates $20$ times in total, with participation every $19$ steps. These numbers were chosen for memory constraints and not optimized. This cyclic participation corresponds to a $(k,b)$-participation scheme with $k=20$ and $b=19$ in the matrix factorization literature (Choquette-Choo et al., 2023c).
@@
We also evaluate Mafalda-SGD on the federated EMNIST (FEMNIST) image classification task from the LEAF benchmark (Caldas et al., 2018), predicting handwritten characters across users. This setting is highly non-iid and realistic for decentralized learning. To ensure enough data is present, each node is given $10$ unique handwriting styles as training data. We use a small convolutional network comprising two convolutional layers (16 and 32 filters, $3\times 3$ kernels) with ReLU and $2\times 2$ max-pooling, followed by group normalization and a fully-connected layer for a softmax output trained with cross-entropy. Training uses a cyclic participation scheme ($(k,b)=(20,16)$).
Figure 4: FEMNIST classification results on the Facebook Ego graph. Left: test accuracy evolution for a fixed privacy budget. Right: privacy-utility tradeoff between the final test accuracy as a function of the LDP privacy budget. Error bands show 95% confidence intervals.
We compare Mafalda-SGD to baselines under local DP accounting. Results (averaged over 20 runs) are reported in Figure 4. Across all privacy budgets, Mafalda-SGD consistently achieves higher test accuracy than the private baselines, with the largest relative gains at high privacy levels where correlated noise preserves utility while satisfying LDP.
@@ Acknowledgments
This work was supported by grant ANR-20-CE23-0015 (Project PRIDE) and the ANR 22-PECY-0002 IPOP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR and by the Austrian Science Fund (FWF) [10.55776/COE12]. This research was supported in part by the Groupe La Poste, sponsor of the Inria Foundation, in the framework of the FedMalin Inria Challenge. This project was provided with computing AI and storage resources by GENCI at IDRIS thanks to the grant 2025-AD011015352R1 on the supercomputer Jean Zay’s V100 partition. Experiments presented in this paper were also carried out using the Grid’5000 testbed, supported by a scientific interest group hosted by Inria and including CNRS, RENATER and several Universities as well as other organizations (see https://www.grid5000.fr). The authors thank Quino for his comic strip Mafalda which inspired our new algorithm name.
Reproducibility
We provide an open-source implementation of the code used in this work, publicly available at https://github.com/dimiarbre/MFDL.